OpenStudy (anonymous):

find the product of (x-3)^2

OpenStudy (jojo4eva):

x^2-6x+9

OpenStudy (anonymous):

Herro Baka *_*

OpenStudy (jojo4eva):

@bangbang559 any questions

OpenStudy (shamil98):

(x-3)(x-3) Use FOIL. or the identity: (x-a)^2 = (x-a)(x-a) = x^2 -a -a + a^2

OpenStudy (anonymous):

@jojo4eva how did you find the answer?

OpenStudy (shamil98):

x^2 - ax - ax + a^2 * my bad.

OpenStudy (anonymous):

@shamil98 your answer is confusing me

OpenStudy (jojo4eva):

what I did was utilize the foil method so (x-3)^2 is the same thing as (x-3)(x-3) then I multiplied eachterm in the first set b and so x^2-3x-3x+9

OpenStudy (anonymous):

okay thanks :)